OpenStudy (cutiecomittee123):

TRIG help Help me to convert these equations from polar to rectangular theta=pi/3 r=5csctheta r^2=sintheta

OpenStudy (danjs):

Polar coordinates are (distance from origin, angle from +x) Rectangular are, (distance from origin X, Distance from origin Y) What do you know about changing from one to the other?

OpenStudy (cutiecomittee123):

Polar to rectangular equation is x=rcostheta y=rsintheta

OpenStudy (danjs):

yeah, using a right triangle you can see those here, with sin and cosine 

OpenStudy (danjs):

(x , y) = (r*cos(theta) , r*sin(theata) )

OpenStudy (cutiecomittee123):

okay

OpenStudy (danjs):

Using the Pythagorean Theorem, you can see [r*cos(theta) ]^2 + [r*sin(theata) ]^2 = r^2 x^2 + y^2 = r^2

OpenStudy (cutiecomittee123):

okay got it :)

OpenStudy (danjs):

and with the tangent of theta, you see tan(theta) = y / x that is really everything

OpenStudy (cutiecomittee123):

okay s if we have r=5csctheta, how d those theorums apply?

OpenStudy (danjs):

r = 5*csc(a) remember csc(a) = 1/sin(a) \[r = 5*\frac{ 1 }{ \sin(\theta) }\]

OpenStudy (danjs):

multiply both sides by sin(a) r*sin(a) = 5 that is the y coordinate (r*sin(a)) y = 5

OpenStudy (danjs):

So the Polar equation \[r = 5*\csc(\theta)\] represents a horizontal line at y=5 in rectangular coordinates

OpenStudy (danjs):

check it, i havent done these in awhile

OpenStudy (danjs):

You see they gave you a value for r, theta is not defined and is still theta, the point is \[(r , \theta) = (5*\csc(\theta) ~,~ \theta)\] You rearrange r=5csc(a), to r*sin(a) = 5, remember y=r*sin(a) y = 5

OpenStudy (danjs):

\[(x , y) = (x , 5)\] horizontal line, y=5, all points (x,5)